OpenStudy (bekkah323):

factor completely: 16y^2+40xy+25x^2

OpenStudy (anonymous):

16y^(2)+40xy+25x^(2) For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (400) that add up to b (40). a=16, b=40, c=25 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (400) that add up to b (40).In this problem (5)/(4)*(5)/(4)=(25)/(16) (which is (c)/(a)) and (5)/(4)+(5)/(4)=(40)/(16) (which is (b)/(a)) , so insert (5)/(4) as the right hand term of one factor and (5)/(4) as the right-hand term of the other factor. (y+(5)/(4)*x)(y+(5)/(4)*x) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (4y+5x)(4y+5x) Combine the two common factors of (4y+5x) by adding the exponents. (4y+5x)^(2)

OpenStudy (bekkah323):

isn't there an equation for this so its shorter?

OpenStudy (anonymous):

(4y+5x)^(2)

OpenStudy (bekkah323):

sorry, i actually meant something like how a^2-b^2= (a+b)(a-b) but for this problem?

OpenStudy (bekkah323):

if that makes sense

OpenStudy (bekkah323):

i know one half of it goes a^2+2ab+b^2 if that applies.